# Understanding and Mitigating the Impacts of Differentially Private Census Data on State Level Redistricting

### v1

@@ 5.1.2 Offsets for state legislative redistricting
We perform a similar analysis for 93 state legislative geographies. For each geography, we measure the discrepancy at $\tau=5\%$ with each of twenty offsets $\Delta$ between 0% and 1% (generating an ensemble of 10,000 plans for each offset). For each geography we record the no-offset discrepancy ($\Delta=0\%$) and the critical offset (where the discrepancy first falls below 2%). Table 1 gives a coarse summary of the results. (See Table 4 in Appendix A for much more detailed results.) Observe that the no-offset discrepancy rates vary widely; in the worst case (Connecticut State House of Representatives) 83% of plans exceeded 5% population deviation under SWAP. But offsets provide a powerful mitigation: for most geographies, $\Delta\leq 0.2\%$ suffices to bring the fraction of problematic plans below 2%.
We do not know why different geographies exhibit different discrepancies and critical offsets, although district size relative to census geographies appears to play an important role. Qualitatively, we find that when districts are many times larger than census tracts, the no-offset discrepancy rate is low; when they are about the same size, it is high. In particular, in state lower house geographies where districts were at least 20 times the size of tracts, the critical offset is 0.1% or less. For the four geographies with fewer than twice as many tracts as districts, the critical offsets are greater than 0.5%. More generally, there is a moderate negative rank correlation between the no-offset discrepancies at $\tau=5\%$ and the ratio of the number of tracts to number of districts (equivalently, the ratio of the average district population to average tract population). Spearman’s rank correlation coefficient is $\rho=-0.645$, with 95% confidence interval (-0.789, -0.434). There is also a strong positive rank correlation between the critical offset and the same ratio ($\rho=0.840$; 95% confidence interval (0.725, 0.909)). We found no evidence of rank correlations with the number of districts in a plan nor the average district population.
Min 25th 50th 75th Max
Discrepancy rate 0.003 0.029 0.090 0.264 0.826
Critical offset (%) 0 0.1 0.1 0.2 1.0
@@ 5.2 Why offsets work: most deviation isn’t from disclosure avoidance
Plotting these two terms side-by-side clarifies their relative import. Figure 2 plots histograms of signed $\mathsf{dev}_{\mathsf{demo}}(D)$ and $\mathsf{err}_{\mathsf{das}}(D)$ for all the LA state senate districts in our $\tau=5\%$ ensembles. Note the different scales on the horizontal axes. The signed $\mathsf{dev}_{\mathsf{demo}}(D)$ is roughly uniform across the acceptable range $[-5\%,5\%]$. The DAS noise $\mathsf{err}_{\mathsf{das}}(D)$ is highly concentrated, with mean 0 and standard deviation 0.060%. Only a small fraction of districts have $\mathsf{dev}_{\mathsf{demo}}(D)$ close enough to $5\%$ for the DAS noise to matter. Using an offset moves $\mathsf{dev}_{\mathsf{demo}}(D)$ away from the 5% threshold, further lowering the number of districts where the DAS noise matters.
@@
Figure 2: The two components of population deviation in districts. Histograms of apparent population deviation in DEMO ($\mathsf{dev}_{\mathsf{demo}}(D)=(\mathsf{pop}_{\mathsf{demo}}(D)-\overline{p})(\overline{p})$) and the additional error from disclosure avoidance ($\mathsf{err}_{\mathsf{das}}(D)=(\mathsf{pop}_{\mathsf{swap}}(D)-\mathsf{pop}_{\mathsf{demo}}(D))/\overline{p}$) for each unique district included in an ensemble of Louisiana state senate plans sampled with a 5% population tolerance on the DEMO data. Note the very different scales on the horizontal axis. Together, these terms make up the population of a state-legislative district’s population deviation. The dashed vertical lines behind the left histogram represent the maximum and minimum error values observed in the right histogram.
@@
Figure 2 suggests a very simple probabilistic model for population deviations under SWAP. The deviation of a plan is sampled as the maximum over district-level deviations: $\mathsf{dev}_{\mathsf{swap}}(\mathcal{P})=\max_{i}\mathsf{dev}_{\mathsf{swap}}(D_{i})$. Each $\mathsf{dev}_{\mathsf{swap}}(D_{i})$ is independently sampled as $|X+E|$, where $X$ is uniform over $[-(\tau-\Delta),\tau-\Delta]$ and $E\sim\mathrm{N}(\mu,\sigma^{2})$. The parameters $\mu=0.0\%$ and $\sigma=0.060\%$ are the empirical mean and standard deviation of $\mathsf{err}_{\mathsf{das}}(D)$ across districts in our LA state senate ensembles.
@@
The dashed lines in Figure 1 show the fraction of 100,000 values of $\mathsf{dev}_{\mathsf{swap}}(\mathcal{P})$ sampled as above that exceeded $\tau$, for each $\tau$ and $\Delta$. Qualitatively, the model closely approximates the empirical data for the LA senate, suggesting this is a useful mental model for understanding the effect of offsets. (Note that this is meant only as a simple model of the phenomenon; in particular, the observed values of $\mathsf{err}_{\mathsf{das}}(D)$ are not normally distributed.)
5.2.1 A closer look at errors from disclosure avoidance
@@
To understand the effect of the 2020 DAS on OPOV questions, it appears that we must understand $|\mathsf{err}_{\mathsf{das}}(D)|$. That is, the magnitude of population error in a district $D$ due to disclosure avoidance as a fraction of the ideal district population. This is perhaps unsurprising. We briefly discuss this error in enacted districts; as compared to other sources of error, and in relation to racial biases in the 2020 DAS.
@@
Real redistricting plans are drawn by people, not sampled from ensembles. We do not know how offsets fare in a real redistricting setting. But the distribution of $|\mathsf{err}_{\mathsf{das}}(D)|$ on districts created by the political process can give us some initial idea. Using DEMO and SWAP, we compute $|\mathsf{err}_{\mathsf{das}}(D)|$ for every congressional and state legislative district enacted after the 2010 Decennial Census (rather than for algorithmic ensembles of districts). We group the districts by ideal population, which doesn’t depend on the redistricting plan. Table 2 reports the maximum, 98th, and 90th percentile values in each group. For the overwhelming majority of geographies—including all with ideal population above $32,000$—the relative error magnitude from disclosure avoidance was well under 1%. A Rhode Island State House of Representatives district saw the largest observed relative error: 2.14%, corresponding to just 301 people.
A redistricter can use these numbers to guide the use of offsets where population deviation under the CEF is a concern, though more research would be warranted. If the ideal district population is $\overline{p}=200,000$, an offset of $\Delta=0.1\%$ might be appropriate; if $\overline{p}=20,000$, an offset of $\Delta=0.63\%$ appears more appropriate.
Ideal population $\overline{p}$ $<$8k 8-16k 16-32k 32-64k 64-128k 128-256k 256-512k $\geq$512k
@@
Table 2: Error from disclosure avoidance in legislative districts enacted after the release of the 2010 census redistricting data, as a fraction of ideal population: $|\mathsf{err}_{\mathsf{das}}(D)|=|\mathsf{pop}_{\mathsf{demo}}(D)-\mathsf{pop}_{\mathsf{swap}}(D)|/\overline{p}$. Districts are grouped by ideal population.
@@
As a point of comparison for the population deviations so far, we analyze district population imbalances arising from population shifts in the 10 years between decennial censuses. This is only one acknowledged source of population deviation among districts, and ignores systematic undercounts and overcounts that differ by race and ethnicity [30]. Using data from IPUMS that standardizes congressional district geographies between decennial censuses, we compare total populations of congressional districts of the 110th-112th Congresses in both the 2000 and 2010 censuses [31]. [^8] Because Census geographies (i.e. block boundaries) are modified every 10 years, districts drawn using 2000 census data may split 2010 census blocks, complicating comparisons between years. When measured with the 2000 data used to draw the districts, the deviations are relatively small. Of states with more than one Congressional district, the average deviation is just 613 people ($\sim$0.1of the ideal population), while the maximum is 6,698 ($\sim$1of the ideal). By the 2010 Census, the average increased to 137,974 people ($\sim$19of the ideal), while the maximum increased to 340,948 people (Texas, $\sim$43of the ideal).
@@
Prior work found that, in precincts, TopDown’s noise is correlated with racial/ethnic homogeneity [6, 32]. [^9] We use the term precinct instead of the more correct “voting district” or VTD to avoid confusion with the much larger state and Congressional districts that are the main subject of this paper [33]. Precincts with higher Herfindahl–Hirschman index—a measure of racial and ethnic homogeneity—tend to have their populations inflated; precincts with lower index, deflated [6]. This motivates the following question: Does the distribution of district-level errors in our ensembles vary by the racial and ethnic makeup of the districts? If so, it might systematically shift voting power among groups even if district populations are still within the acceptable bounds. Fortunately, we do not observe any meaningful effect. Using of the 1,626,525 unique districts we sampled for the Georgia state house, a geography for which small changes in voting power of Black residents could have a meaningful impact on the makeup of the legislature, we find that the Pearson correlation between per-district population error and Herfindahl-Hirschman index is minute: $r=0.029$, 95% CI $(0.028,0.031)$.
@@ 6 Counting majority-minority districts with noise
To challenge an enacted electoral map for racial vote dilution, plaintiffs must meet the Gingles 1 precondition. Gingles 1 requires showing that one can draw more majority-minority districts (MMDs) than exist in the enacted map being challenged. This is usually done by submitting expert reports with one or more such illustrative plans. As with population deviation, the apparent number of MMDs in a plan may differ by dataset. Absent clear law about which dataset is the appropriate benchmark, Alabama worried that this could make gerrymandering lawsuits more likely, while the ACLU worried this could make those lawsuits less successful (Section 1).
This section asks whether one can show that there are plans with more MMDs than an enacted plan, according to the SWAP dataset (standing in for the 2020 CEF) but using only the DEMO dataset (standing in for the 2020 Redistricting Data). We specifically focus on majority-Black districts. As is typical in VRA litigation, majorities are measured in the voting age population rather than the total population (Georgia v. Ashcroft (2003)).
@@
We generate ensembles of plans using a technique called short-burst optimization, which is designed to output plans with many majority-Black MMDs. This is motivated by the observation that Gingles 1 incentivizes plaintiffs to maximize the number of MMDs in the illustrative plan. (See 6.2.1 for a comparison to the base ensemble.) It is generally intractable to determine the maximum possible number of MMDs in a plan at the scale of a US state. Therefore, we use the optimized chains as a proxy for a redistricter who is trying to draw a plan with a near-maximal number of majority-minority districts. While the human and the algorithm differ in their methods for selecting plans, this methodology sheds light on the qualities of plans that are drawn from a nearly-maximal distribution.
@@
We examine the MMD discrepancy of plans generated using the DEMO dataset: $\mathsf{MMD}_{\mathsf{demo}}(\mathcal{P})-\mathsf{MMD}_{\mathsf{swap}}(\mathcal{P})$. We measure the mean MMD discrepancy and the non-zero discrepancy rate—the fraction of plans with non-zero MMD discrepancy. Section 6.1 shows that MMD discrepancies can be significant, using the GA state house as a case study. It argues that, even so, the Gingles 1 test can continue to be used as is. Section 6.2 explores the phenomenon of MMD discrepancies more deeply.
6.1 Gingles 1 in the face of MMD discrepancy

### v2

@@ 5.1.2 Offsets for State Legislative Redistricting
We perform a similar analysis for 93 state legislative geographies. For each geography, we measure the discrepancy at $\tau=5\%$ as we increase $\Delta$ in steps of 0.05% (generating an ensemble of 100,000 plans for each offset). For each geography we record the no-offset discrepancy ($\Delta=0\%$) and the critical offset (where the discrepancy first falls to 0%).
Table 1 gives a coarse summary of the results (Table 4 in Appendix A gives much more detail). Observe that the no-offset discrepancy rates vary widely. In the worst case (Kansas State House of Representatives) 96% of plans exceeded 5% population deviation under SWAP. But offsets provide a powerful mitigation: for most geographies, $\Delta\leq 0.75\%$ suffices to eliminate plans that exceed a 5% population tolerance from our ensembles.
Min $25^{th}$ $50^{th}$ $75^{th}$ Max
Discrepancy Rate 4.73% 22.65% 40.59% 66.13% 100.00%
Critical Offset 0.10% 0.28% 0.44% 0.74% 2.18%
@@
We do not know why different geographies exhibit different discrepancies and critical offsets, although the size of districts appears to play an important role. In Figure 2, we examine the standard deviations of the observed district population errors in our state lower and upper legislative district ensembles. Notably, we observe a strong positive correlation between the logarithm of average district population and error standard deviation in both state lower houses ($r=0.87$) and state upper houses ($r=0.83$). We also observe weaker correlations between the logarithm of total state population and error standard deviation (lower houses: $r=0.64$, upper houses: $r=0.82$) and between the logarithm of the average number of census blocks per district and error standard deviation (lower houses: $r=0.59$, upper houses: $r=0.66$).
Figure 2: Standard deviation of population error across state legislative geographies. Standard deviations of the population error ($\mathsf{pop}_{\mathsf{demo}}(D)-\mathsf{pop}_{\mathsf{swap}}(D)$) over districts in ensembles generated for state upper and lower legislative districting plans. Each point represents a state legislative geography. The left column plots the logarithm of each state’s population on the horizontal axis. The center column plots the logarithm of each legislative geography’s average district population on the horizontal axis. The right column plots the logarithm of the total number of blocks in each state divided by the number of districts in each respective legislative geography on the horizontal axis.
We ran additional experiments on the state of Louisiana to understand the link between district size and discrepancy rate. For each value of $\Delta$ we examined in Figure 1, we measure the discrepancy rate at $\tau=5\%$ for plans with different numbers of districts, ranging from 5 to 160. This corresponds to average district populations between 28,334 and 906,674 people. [^10] For comparison: Louisiana Congressional districts contain about 766,000 people; New Orleans City Council districts contain about 72,500 people. Results are shown in Figure 3. We observe that for all values of $\tau-\Delta$, the discrepancy rate is monotonically decreasing as the ideal district size increases.
Figure 3: Discrepancies by district size in Louisiana. Measuring the discrepancy rate of Louisiana redistricting plans over various district sizes and population tolerance offsets. Each point represents an ensemble that was generated using a population tolerance threshold of $5\%-\Delta$ on DEMO (for several values of $\Delta$) and evaluated with a $\tau=5\%$ threshold on SWAP. The ideal district populations correspond to plans with 5, 10, 20, 40, 80, 120, and 160 plans. The vertical red line corresponds with the actual district size of the Louisiana State Senate.
@@ 5.2 Why Offsets Work: Most Deviation isn’t from Disclosure Avoidance
Plotting these two terms side-by-side clarifies their relative import. Figure 4 plots histograms of signed $\mathsf{dev}_{\mathsf{demo}}(D)$ and $\mathsf{err}_{\mathsf{das}}(D)$ for all the LA state senate districts in our $\tau=5\%$ ensembles. Note the different scales on the horizontal axes. The signed $\mathsf{dev}_{\mathsf{demo}}(D)$ is roughly uniform across the acceptable range $5\%,5\%]$. The DAS noise $\mathsf{err}_{\mathsf{das}}(D)$ is highly concentrated, with mean 0.0% and standard deviation 0.080%. Only a small fraction of districts have $\mathsf{dev}_{\mathsf{demo}}(D)$ close enough to $5\%$ for the DAS noise to matter. Using an offset moves $\mathsf{dev}_{\mathsf{demo}}(D)$ away from the 5% threshold, further lowering the number of districts where the DAS noise matters.
@@
Figure 4: The two components of population deviation in districts. Histograms of apparent population deviation in DEMO ($\mathsf{dev}_{\mathsf{demo}}(D)=(\mathsf{pop}_{\mathsf{demo}}(D)-\overline{p}/\overline{p}$) and the additional error from disclosure avoidance ($\mathsf{err}_{\mathsf{das}}(D)=(\mathsf{pop}_{\mathsf{swap}}(D)-\mathsf{pop}_{\mathsf{demo}}(D))/\overline{p}$) for each unique district included in an ensemble of Louisiana state senate plans sampled with a 5% population tolerance on the DEMO data. Note the very different scales on the horizontal axis. Together, these terms make up the population of a state legislative district’s population deviation. The dashed vertical lines behind the left histogram represent the maximum and minimum error values observed in the right histogram.
@@
Figure 4 suggests a very simple probabilistic model for population deviations under SWAP. The deviation of a plan is sampled as the maximum over district-level deviations: $\mathsf{dev}_{\mathsf{swap}}(\mathcal{P})=\max_{i}\mathsf{dev}_{\mathsf{swap}}(D_{i})$. Each $\mathsf{dev}_{\mathsf{swap}}(D_{i})$ is independently sampled as $|X+E|$, where $X$ is uniform over $(\tau-\Delta),\tau-\Delta]$ and $E\sim\mathrm{N}(\mu,\sigma^{2})$. The parameters $\mu=0.0\%$ and $\sigma=0.080\%$ are the empirical mean and standard deviation of $\mathsf{err}_{\mathsf{das}}(D)$ across districts in our LA state senate ensembles.
@@
The dashed lines in Figure 1 show the fraction of 100,000 values of $\mathsf{dev}_{\mathsf{swap}}(\mathcal{P})$ sampled as above that exceeded $\tau$, for each $\tau$ and $\Delta$. Qualitatively, the model closely approximates the empirical data for the LA senate, suggesting this is a useful mental model for understanding the effect of offsets.
Note that this is meant only as a simple, usable model of the phenomenon. In particular, the observed values of $\mathsf{err}_{\mathsf{das}}(D)$ do not appear normally distributed, failing the parametric bootstrap goodness of fit test implemented in SciPy [46].)
5.3 A Closer Look at Errors from Disclosure Avoidance
@@
To understand the effect of the 2020 DAS on OPOV questions, it appears that we must understand $|\mathsf{err}_{\mathsf{das}}(D)|$: the magnitude of population error in a district $D$ due to disclosure avoidance, as a fraction of the ideal district population. This is perhaps unsurprising.
We briefly discuss this error in enacted districts, as compared to other sources of error, and in relation to racial biases in the 2020 DAS.
@@
Real redistricting plans are drawn by people, not sampled from ensembles. We do not know how offsets fare in a real redistricting setting. But the distribution of $|\mathsf{err}_{\mathsf{das}}(D)|$ on districts created by the political process can give us some initial idea. Using DEMO and SWAP, we compute $|\mathsf{err}_{\mathsf{das}}(D)|$ for every congressional and state legislative district enacted after the 2010 Decennial Census (rather than for algorithmic ensembles of districts). We group the districts by ideal population, which doesn’t depend on the redistricting plan. Table 2 reports the maximum, 98th, and 90th percentile values in each group. For the overwhelming majority of geographies—including all with ideal population above 32,00032,000—the relative error magnitude from disclosure avoidance was well under 1%. A Rhode Island State House of Representatives district saw the largest observed relative error: 2.14%, corresponding to just 301 people.
A redistricter can use these numbers to guide the use of offsets where population deviation under the CEF is a concern, though more research would be warranted. If the ideal district population is $\overline{p}=200,000$, an offset of $\Delta=0.51\%$ might be a conservative choice; if $\overline{p}=20,000$, an offset of $\Delta=1.37\%$ appears more appropriate.
Ideal population $\overline{p}$ <<8k 8-16k 16-32k 32-64k 64-128k 128-256k 256-512k $\geq$512k
@@
Table 2: Error from disclosure avoidance in state legislative districts and Congressional districts enacted after the release of the 2010 census redistricting data, as a fraction of ideal population: $|\mathsf{err}_{\mathsf{das}}(D)|=|\mathsf{pop}_{\mathsf{demo}}(D)-\mathsf{pop}_{\mathsf{swap}}(D)|/\overline{p}$. Districts are grouped by ideal population.
@@
As a point of comparison for the population deviations so far, we analyze district population imbalances arising from population shifts in the 10 years between decennial censuses. This is only one acknowledged source of population deviation among districts, and ignores systematic undercounts and overcounts that differ by race and ethnicity [44]. Using data from IPUMS that standardizes congressional district geographies between decennial censuses, we compare total populations of congressional districts of the 110th-112th Congresses in both the 2000 and 2010 censuses [30]. [^12] Because Census geographies (i.e. block boundaries) are modified every 10 years, districts drawn using 2000 census data may split 2010 census blocks, complicating comparisons between years. When measured with the 2000 data used to draw the districts, the deviations are relatively small. Of states with more than one Congressional district, the average deviation is just 613 people ($\sim$0.1% of the ideal population), while the maximum is 6,698 ($\sim$1% of the ideal). By the 2010 Census, the average increased to 137,974 people ($\sim$19% of the ideal), while the maximum increased to 340,948 people (Texas, $\sim$43% of the ideal).
@@
Prior work found that noise from the 2020 DAS is correlated with racial/ethnic homogeneity [24, 35]. Specifically, precincts with higher Herfindahl–Hirschman index—a measure of racial and ethnic homogeneity—tend to have their populations inflated, and vice versa. [24]. This motivates the following question: Does the distribution of district-level errors in our ensembles vary by the racial and ethnic makeup of the districts? If so, it might systematically shift voting power among groups even if district populations are still within the acceptable bounds. Fortunately, we do not observe any meaningful effect. For the 876,800 unique districts we sampled for the Georgia state house, a geography for which small changes in voting power of Black residents could have a meaningful impact on the makeup of the legislature, we find that the Pearson correlation between per-district population error and Herfindahl-Hirschman index is minute: $r=0.07$, 95% CI $(0.068,0.072)$.
@@ 6 Counting Majority-Minority Districts with Noise
To challenge an enacted electoral map for racial vote dilution, plaintiffs must meet the Gingles 1 precondition. Gingles 1 requires showing that one can draw more majority-minority districts (MMDs) than exist in the enacted map being challenged. This is usually done by submitting expert reports with one or more such illustrative plans.
This section investigates how perturbations from disclosure avoidance interact with Gingles 1. We consider a data user’s ability to use the public DEMO dataset (after disclosure avoidance) to produce illustrative plans with many majority-minority districts under the confidential CEF dataset (before disclosure avoidance). Notice that these illustrative plans may be quite different than those produced in the normal course of redistricting: they must have more majority-minority districts to satisfy Gingles 1. For this reason, we also investigate the additional impact of optimizing the number of majority-minority districts using a technique called short-burt optimization [8] described below.
[…]
This section specifically focuses on majority-Black districts and uses MMD to mean a district where a majority of the voting age population is Black, excluding those who select multiple races. As before, we restrict our attention to state legislative redistricting.
6.1 Short-Burst Optimization
@@
Gingles 1 incentivizes plaintiffs to maximize the number of MMDs in the illustrative plan, shifting the distribution of plans. To study the effect of this shift, we generate ensembles of plans using a technique called short-burst optimization, which is designed to output plans with many majority-Black MMDs [8]. It is generally intractable to determine the maximum possible number of MMDs in a plan at the scale of a US state. Therefore, we use the optimized chains as a proxy for a redistricter who is trying to draw a plan with a near-maximal number of majority-minority districts.
We use short-burst optimization in a manner suggested to us by Duchin (personal communication, August 2025). We run 100 chains (500 for Georgia) with short burst optimization for 100,000 steps each in parallel. From each chain, we randomly subsample 10 plans from among those with highest MMD count in DEMO, resulting in 1,000-plan ensembles (5,000 for Georgia).
This approach is somewhat heuristic, but it allows us to observe the effects of incentivizing maps with certain characteristics, as Gingles 1 does. The ensembles can also be seen as containing plans that might have been selected as a Gingles 1 illustrative plan by a redistricter who used short-burst optimization to find maps with many MMDs. While the human and the algorithm differ in their methods for selecting plans, this methodology sheds light on the qualities of plans that are drawn from a nearly-maximal distribution.
This approach is computationally intensive, requiring hundreds of CPU-hours to generate an ensembles of 5,000 plans for the state of Georgia. We ran fewer chains for other states in order to allow us to complete our experiments in a reasonable timeframe. We observed that analyzing the first 1,000 plans of the Georgia ensemble yielded similar results to the full 5,000 plan ensemble.
6.2 Measuring MMD Discrepancy
We begin with a case study of Georgia state house plans, then extend the analysis to 50 state legislative geographies. The Georgia house serves as a good illustrative geography due to its significant Black population and large legislature.
@@ 6.2 Measuring MMD Discrepancy
We examine the MMD discrepancy of plans generated using the DEMO dataset: $\mathsf{MMD}_{\mathsf{demo}}(\mathcal{P})-\mathsf{MMD}_{\mathsf{swap}}(\mathcal{P})$. We measure the mean MMD discrepancy and the non-zero discrepancy rate—the fraction of plans with non-zero MMD discrepancy.
6.2.1 Georgia State House Plans